### v1

@@ Abstract
Dataset curation has become a basis for strong large language model (LLM) performance. While various rule-based filtering heuristics exist for English and multilingual datasets, model-based filtering techniques have primarily focused on English. To address the disparity stemming from limited research on non-English languages, we propose a model-based filtering framework for multilingual datasets that aims to identify a diverse set of structured and knowledge-rich samples. Our approach emphasizes transparency, simplicity, and efficiency, leveraging Transformer- and FastText-based classifiers to ensure the broad accessibility of our technique and data. We conduct comprehensive ablation studies on the FineWeb-2 web crawl dataset across diverse language families, scripts, and resource availability to demonstrate the effectiveness of our method. Training a 1B-parameter Llama model for 70B and 119B tokens, our approach can match the baseline MMLU score with as little as 15% of the training tokens, while also improving across other benchmarks. These findings provide strong evidence for the generalizability of our approach to other languages. As a result, we extend our framework to 20 languages for which we release the refined pretraining datasets.
@@ 1 Introduction
Deduplication and heuristic-based dataset cleaning have become standard practices in data curation (Rae et al., 2021; Raffel et al., 2020; De Gibert et al., 2024). These quality filters are often complemented by additional filters, such as the removal of personally identifiable information (PII) (Penedo et al., 2024a) or model-based toxicity filtering (Soldaini et al., 2024). Recently, model-based filtering has also emerged as a promising method for quality filtering. The release of FineWeb-Edu (Penedo et al., 2024a) demonstrated that pretraining on just 10% of the tokens (38B) from an English dataset filtered using a model-based approach can achieve performance comparable to models trained on 350B tokens of unfiltered data. Moreover, when trained on equivalent amounts of data, this model largely outperforms the baseline. Concurrently, the release of DCLM (Li et al., 2024b) showed that competitive performance can be achieved using a simple and efficient model-based approach, namely a FastText (Joulin et al., 2017) classifier trained on a carefully selected training dataset.
@@
To achieve this, we leverage embedding-based classification models. Firstly, we adopt the FastText quality filtering approach from DCLM to develop a unified framework for multilingual datasets that span diverse language families, scripts, and resource availability, focusing on Chinese, German, French, Arabic, and Danish as representative languages for our experiments. Additionally, we extend this embedding-based approach by incorporating Transformer (Vaswani et al., 2023) embeddings, specifically XLM-RoBERTa (Conneau et al., 2020), for filtering. We compare the performance between baseline FineWeb-2 data and our best FastText and Transformer embedding-based approaches in Figure 1.
@@
We propose a transparent, simple, and unified framework for multilingual model-based filtering at web scale, enabling data curation across diverse language families, scripts and resource availability.
We present comprehensive per-language ablation studies of embedding-based multilingual quality filtering on top of the FineWeb-2 dataset (Penedo et al., 2024c), achieving performance comparable to the baseline while using as little as 15% of the tokens. We additionally analyze the impact of dataset contamination and multilingual LLM training.
We evaluate the impact of different training datasets for data selection classifiers on the downstream performance of LLMs.
We release the refined pretraining dataset222huggingface.co/epfml covering 20 languages333Russian, Chinese, German, Japanese, Spanish, French, Italian, Portuguese, Polish, Dutch, Indonesian, Turkish, Czech, Vietnamese, Swedish, Persian, Arabic, Greek, Danish, Hungarian, filtered using our proposed framework, along with the codebase444github.com/epfml/fineweb2-hq, to advance multilingual language modeling.
@@ 2 Related Work
Data Curation. In order to pretrain LLMs on a large amount of diverse texts, Common Crawl555commoncrawl.org is often used as the base dataset. However, early works already observed that performing quality filtering on Common Crawl is crucial for model performance (Brown et al., 2020). There exist various data curation approaches, such as deduplication (Lee et al., 2022), PII removal (Subramani et al., 2023), or toxicity filtering (Arnett et al., 2024). Another important aspect is quality filtering of the documents. For this, the definition of quality is an important aspect. A common approach is to use heuristics to remove documents outside of the target distribution, such as filtering based on average word length, existence of punctuation, or document length (Rae et al., 2021; Raffel et al., 2020). Another approach is to define model-based filters, where research has focused on perplexity measure of the text (Wenzek et al., 2019) or focused on educational (Penedo et al., 2024a) and conversational documents (Li et al., 2024b). In this work, we build upon previous curated datasets based on heuristic filtering, specifically Finweb-2 (Penedo et al., 2024c), and focus on model-based filtering for structured and knowledge-rich documents relying on textual embedding representation.
@@
Curated English datasets. One of the early curated datasets was C4 (Raffel et al., 2020), followed by MassiveText (Rae et al., 2021). RefinedWeb (Penedo et al., 2023) was an important step forward, demonstrating that filtered web data can outperform selected high-quality data sources. While these datasets have not been made fully publicly available, their filtering techniques have been expanded upon in recent fully public datasets, such as Dolma (Soldaini et al., 2024), FineWeb, and FineWeb-Edu (Penedo et al., 2024a). While FineWeb primarily relies on filter heuristics for data quality, Dolma adopts model perplexity filtering. FineWeb-Edu takes model-based filtering a step further and relies on LLM-based quality assessment. Similarly, a concurrent work, DCLM, has achieved competitive performance using FastText (Joulin et al., 2017) classifier trained on a carefully selected training dataset. In this work we adapt and extend this approach to the multilingual context.
@@
Curated Multilingual Datasets. Analogously to the English datasets, there have been efforts in the multilingual space. An influential work has been CCNet (Wenzek et al., 2019), whose language identification and model perplexity filter for data quality has been re-used in later datasets. Again, while CCNet was not published directly, but rather provided the tools for data cleaning, RedPajama (Together Computer, 2023) is a prominent multilingual dataset relying on these filtering techniques. While RedPajama offers data in 5 European languages, other datasets, such as OSCAR (Ortiz Suárez et al., 2019; Abadji et al., 2021; 2022), mC4 (Xue et al., 2021), ROOTS (Laurençon et al., 2022), MADLAD-400 (Kudugunta et al., 2023), CulturaX (Nguyen et al., 2023), and HPLT (de Gibert et al., 2024), focus on expanding beyond, spanning a variety of language families and scripts. While they offer refined datasets for hundreds of languages, FineWeb-2 (Penedo et al., 2024c) pushes the limit to thousands of languages and further improves the performance. Our work also focuses on filtering quality samples across various language families and scripts. However, we limit our scope to 20 languages, as the number of documents drops quickly and there is trade-off between retaining a sufficient number of pretraining tokens and ensuring data quality (Muennighoff et al., 2023; Held et al., 2025). In our results, we observe the greatest benefits using stricter data filtering.
@@
Multilingual Embedding Models. Early word embedding models like Word2Vec (Mikolov et al., 2013) and GloVe (Pennington et al., 2014) lacked contextual understanding. FastText (Bojanowski et al., 2017) built upon them and improved performance by incorporating subword information. Transformer (Vaswani et al., 2023) models like BERT (Devlin et al., 2019) and GPT (Radford et al., 2018) then revolutionized the field with context-aware embeddings. Multilingual models like mBERT, XLM (Lample & Conneau, 2019), and XLM-RoBERTa (Conneau et al., 2020) further advanced cross-lingual understanding, with recent open-source LLMs pushing performance even higher (Llama Team, 2024; Mistral AI, 2025). Using such models, documents as well as representative samples can be mapped into a shared embedding space to estimate their similarity. Focusing on transparency, simplicity and efficiency in our work, we use FastText and XLM-RoBERTa for our filtering, and analyze the trade-off between computational complexity and filtering performance.
@@
Multilingual Evaluation. Evaluating LLMs requires diverse benchmarks testing linguistic and cognitive abilities like reading comprehension, reasoning, and knowledge. While English benchmarks like MMLU (Hendrycks et al., 2020) and ARC (Clark et al., 2018) exist, other languages often use translations from English, e.g., XNLI (Conneau et al., 2018) and machine-translated version of MMLU (Lai et al., 2023). However, translations can be problematic, failing to capture cultural nuances or introducing ”translationese” (Romanou et al., 2024). Recent work by Romanou et al. (2024); Singh et al. (2024a) emphasizes the need for culturally sensitive, natively collected benchmarks. Task difficulty and task formulation also impact model performance when trained for shorter durations (Kydlíček et al., 2024). In our work, we follow the recent evaluation tasks selection and methodology by Kydlíček et al. (2024) to assess our model-based filtering approaches across multiple languages.
@@ 3.1 Classifier Training Dataset
Aya Collection. A prompt completion dataset comprising 514M samples covering a wide variety of tasks, generated using instruction-style templates in 101 languages (Singh et al., 2024b).
@@
Include-Base-44. Multiple-choice questions focused on general and regional knowledge, as well as reasoning, extracted from academic and professional exams. Spanning 44 languages, it includes a total of 23K samples (Romanou et al., 2024).
@@
To address this quality difference, we create two Multilingual Knowledge Collection (MKC) configurations:
@@
This allows us to evaluate the trade-off between data quality and scale.
Dataset Creation. For our model-based filtering approaches, our goal is to identify documents from the pretraining dataset that are most similar to our representative samples, with the notion of similarity determined by the specific classifier used. We can measure the similarity to our training dataset directly, for example, by computing the cosine similarity to our training samples in the embedding space. Alternatively, following the approach of Li et al. (2024b), the task can be framed as a binary classification problem, with the representative samples as the positive class. For the negative class, we can simply subsample documents from our pretraining dataset, under the assumption that the majority of these documents are neither well-structured nor knowledge-rich. We use both approaches for our classifiers.
To create the binary classification training dataset, we selected 80K random examples from the training set (MKC or MKC+) as positive samples and 80K random examples from FineWeb-2 as negative samples. For smaller datasets, such as Include-Base-44, the entire dataset was used. The same training dataset was utilized across all model-based filtering approaches, disregarding negative samples when unnecessary. Additionally, we created a training dataset for each language individually to avoid leaking language-specific biases to data of other languages.
@@ 3.2 FastText-based Filtering (FT)
To efficiently process datasets with over 100 million documents (Penedo et al., 2024c), similar to DCLM (Li et al., 2024b), we used a binary FastText classifier (Joulin et al., 2017). This classifier runs on CPU and can be easily deployed across multiple cores, for example using DataTrove (Penedo et al., 2024b).
We trained our FastText classifier on the processed training set using 2-gram features (4-gram for Chinese). Additional details about the training process are given in Appendix A.1. These classifiers were then used to assign scores to all documents in the pretraining dataset. To filter the dataset, we applied a score threshold based on the desired retention percentage of documents. This approach balances dataset size and the predicted quality of the samples.
@@ 3.3 Transformer Embedding-based Filtering
To leverage rich semantic information based on contextual relationships, we utilized the Transformer model embeddings. Specifically, we selected a pretrained XLM-RoBERTa base model (Conneau et al., 2020) due to its support of 100 languages, a relatively small size of approximately 279M parameters, and its transparent training procedure. This choice enabled us to process web-scale data efficiently without being restricted to a single language and to align with our commitment to open science.
@@
Multi-Layer Perceptron (MLP). We trained a single-hidden-layer neural network with a hidden dimension of 256, the ReLU activation function, a dropout rate of 20%, and the sigmoid function on the output. The network was trained for 6 epochs using the AdamW optimizer (Loshchilov, 2017) with a constant learning rate and binary cross-entropy loss. We computed document scores using the output layer of the MLP model, which used XML-RoBERTa document embeddings as input.
@@ 4.1 Experimental Setup
Technical Details. We evaluate 1B-parameter Llama models (Llama Team, 2024) to demonstrate the effectiveness of our model-based filtering approaches. The models are trained on either 70B or 119B tokens, balancing token quality and diversity. The smaller dataset (70B tokens) exposes the model to each token at most once (with a few exceptions where some tokens appear twice). The larger dataset (119B tokens) simulates longer training, resulting in increased token repetition. Training utilizes the HuggingFace Nanotron library (Hugging Face, 2024a) with the AdamW optimizer (Loshchilov, 2017) and a WSD learning rate schedule (Hägele et al., 2024).
@@
To minimize the need for costly hyperparameter tuning, we maintain a consistent setup across all experiments. Specifically, we adopt the DeepSeek scaling law (DeepSeek-AI et al., 2024) with a batch size of 1.6M tokens, learning rate of 0.0008, and 2000 warmup steps. We provide our Nanotron config in Appendix A.2.
@@
As base dataset, we use FineWeb-2 (Penedo et al., 2024c), which has been shown to provide a strong baseline across a variety of languages. Since FineWeb-2 is globally deduplicated, we rehydrate both filtered and unfiltered data using the hyperparameters recommended by Penedo et al. (2024c).
@@
To validate our method on English, we use three datasets: FineWeb (Penedo et al., 2024a) as the baseline, along with FineWeb-Edu (Penedo et al., 2024a) and DCLM (Li et al., 2024b), both of which represent the current state-of-the-art. Tokenization is performed using the multilingual Mistral v3 (Tekken) tokenizer (Mistral AI, 2024). All experiments are conducted using 80 NVIDIA GH200 chips.
@@
Evaluation. Our evaluation prioritizes a diverse range of tasks to ensure the models retain well-rounded capabilities, rather than focusing exclusively on knowledge-based tasks. Specifically, we include tasks covering reading comprehension, general knowledge, natural language understanding, common-sense reasoning, and generative tasks in the target language. To evaluate our approach, we use the HuggingFace LightEval library (Fourrier et al., 2023).
@@
For French, Chinese, and Arabic, we utilize the FineTasks (Kydlíček et al., 2024) multilingual evaluation suite, which is designed to provide meaningful signals even for models trained in the order of 100B tokens. We select analogous tasks for German and Danish. For English, we rely on the SmolLM tasks suite (Hugging Face, 2024b). A complete list of tasks and their evaluation metrics for each language is provided in Appendix C.
@@
Model Selection. We follow the approach used in FineTasks for filter selection, computing a global rank score across individual metrics and languages to determine the optimal approach.

### v2

@@ Abstract
Dataset curation has become a basis for strong large language model (LLM) performance. While various rule-based filtering heuristics exist for English and multilingual datasets, model-based filtering techniques have primarily focused on English. To address the disparity stemming from limited research on non-English languages, we develop a model-based filtering framework for multilingual datasets that aims to identify a diverse set of structured and knowledge-rich samples. Our approach emphasizes transparency, simplicity, and efficiency, leveraging Transformer- and FastText-based classifiers to ensure the broad accessibility of our technique and data. We conduct comprehensive ablation studies on the FineWeb-2 web crawl dataset across diverse language families, scripts, and resource availability to demonstrate the effectiveness of our method. Training a 1B-parameter Llama model for 70B and 119B tokens, our approach can match the baseline MMLU score with as little as 15% of the training tokens, while also improving across other benchmarks and mitigating the curse of multilinguality. These findings provide strong evidence for the generalizability of our approach to other languages. As a result, we extend our framework to 20 languages for which we release the refined pretraining datasets.
@@ 1 Introduction
Deduplication and heuristic-based dataset cleaning have become standard practices in data curation (Rae et al., 2021; Raffel et al., 2020; De Gibert et al., 2024). These quality filters are often complemented by additional filters, such as the removal of personally identifiable information (PII) (Penedo et al., 2024b) or model-based toxicity filtering (Soldaini et al., 2024). Recently, model-based filtering has also emerged as a promising method for quality filtering. The release of FineWeb-Edu (Penedo et al., 2024b) demonstrated that pretraining on just 10% of the tokens (38B) from an English dataset filtered using a model-based approach can achieve performance comparable to models trained on 350B tokens of unfiltered data. Moreover, when trained on equivalent amounts of data, this model largely outperforms the baseline. Concurrently, the release of DataComp-LM (DCLM) (Li et al., 2024b) showed that competitive performance can be achieved using a simple and efficient model-based approach, namely a FastText (Joulin et al., 2017) classifier trained on a carefully selected training dataset.
@@
To achieve this, we leverage embedding-based classification models. Firstly, we adopt the FastText quality filtering approach from DCLM to develop a unified framework for multilingual datasets that span diverse language families, scripts, and resource availability, focusing on Chinese, German, French, Arabic, and Danish as representative languages for our experiments. Additionally, we extend this embedding-based approach by incorporating Transformer (Vaswani et al., 2023) embeddings, specifically XLM-RoBERTa (Conneau et al., 2020), for filtering. Figure 1 shows the clear performance gains of our best FastText and Transformer embedding-based approaches over the state-of-the-art baseline FineWeb-2 data.
@@
We develop a transparent, simple, and unified framework for multilingual model-based filtering at web scale, enabling data curation across diverse language families, scripts and resource availability.
We present comprehensive per-language ablation studies of embedding-based multilingual quality filtering on top of the FineWeb-2 dataset (Penedo et al., 2024c), achieving performance comparable to the baseline while using as little as 15% of the tokens. We additionally analyze the impact of dataset contamination. Lastly, our experiments show that our dataset doesn’t suffer from the curse of multilinguality (Chang et al., 2023).
We evaluate the impact of different data selection classifiers, in particular their training datasets, on the downstream performance of LLMs.
We release the refined pretraining dataset222huggingface.co/datasets/epfml/FineWeb2-HQ covering 20 languages333Russian, Chinese, German, Japanese, Spanish, French, Italian, Portuguese, Polish, Dutch, Indonesian, Turkish, Czech, Vietnamese, Swedish, Persian, Arabic, Greek, Danish, Hungarian (dataset details in Appendix A), filtered using our proposed framework, along with the codebase444github.com/epfml/fineweb2-hq, to advance multilingual language modeling.
@@ 2 Related Work
Data Curation. In order to pretrain LLMs on a large amount of diverse texts, Common Crawl555commoncrawl.org is often used as the base dataset. However, early works already observed that performing data curation on Common Crawl is crucial for model performance (Brown et al., 2020). In fact, data curation is important across NLP tasks (Peter et al., 2023; Finkelstein et al., 2024). Specifically for pretraining data, there exist various data curation approaches, such as deduplication (Lee et al., 2022), PII removal (Subramani et al., 2023), or toxicity filtering (Arnett et al., 2024). Another important aspect is quality filtering of the documents. For this, the definition of quality is an important aspect. A common approach is to use heuristics to remove documents outside of the target distribution, such as filtering based on average word length, existence of punctuation, or document length (Rae et al., 2021; Raffel et al., 2020). Another approach is to define model-based filters, where research has focused on perplexity measure of the text (Wenzek et al., 2019; Marion et al., 2023; Ankner et al., 2024), distributional similarity measures (Li et al., 2024b) and LLM-based quality assessment (Gunasekar et al., 2023; Wettig et al., 2024; Sachdeva et al., 2024; Penedo et al., 2024b). In this work, we build upon previous curated datasets based on heuristic filtering, namely the state-of-the-art dataset FineWeb-2 (Penedo et al., 2024c), and focus on model-based filtering for structured and knowledge-rich documents relying on textual embeddings.
@@
Curated English datasets. One of the early curated datasets was C4 (Raffel et al., 2020), followed by MassiveText (Rae et al., 2021). RefinedWeb (Penedo et al., 2023) was an important step forward, demonstrating that filtered web data can outperform selected high-quality data sources. Although these datasets have not been made fully publicly available, their filtering techniques have been expanded upon in recent fully public datasets, such as Dolma (Soldaini et al., 2024), FineWeb, FineWeb-Edu (Penedo et al., 2024b) and DCLM (Li et al., 2024b). While FineWeb primarily relies on filter heuristics for data quality, Dolma adopts model perplexity filtering. FineWeb-Edu takes model-based filtering a step further and relies on LLM-based quality assessment. DCLM, a concurrent work, has achieved competitive performance using a FastText (Joulin et al., 2017) classifier trained on a carefully selected training dataset. In this work we adapt and extend this approach to the multilingual context.
@@
Curated Multilingual Datasets. Analogously to English datasets, significant work has been done in the multilingual space. For example, CCNet (Wenzek et al., 2019) has been influential, with its language identification and model perplexity filtering for data quality being adopted in subsequent datasets. Similar to earlier English datasets, CCNet was not published directly, but rather provided tools for data cleaning. RedPajama (Together Computer, 2023) is a prominent multilingual dataset relying on these filtering techniques, offering data in 5 European languages. Other datasets, such as OSCAR (Ortiz Suárez et al., 2019; Abadji et al., 2021, 2022), mC4 (Xue et al., 2021), ROOTS (Laurençon et al., 2022), MADLAD-400 (Kudugunta et al., 2023), CulturaX (Nguyen et al., 2023), and HPLT (de Gibert et al., 2024), expanded coverage across a variety of language families and scripts. These datasets offer refined content for hundreds of languages, while FineWeb-2 (Penedo et al., 2024c) pushes the limit to thousands of languages and further improves performance. A concurrent work by Martins et al. (2025) uses translation to train a multilingual quality filter based on the English FineWeb-Edu (Penedo et al., 2024b) scores. Our work similarly focuses on filtering high-quality samples across various language families and scripts. However, we take a different approach: we train a separate classifier for each language from scratch using structured and knowledge-rich representative samples. We limit our scope to 20 languages as the number of documents drops quickly and there is trade-off between retaining a sufficient number of pretraining tokens and ensuring data quality (Muennighoff et al., 2023; Held et al., 2025).
@@
Multilingual Embedding Models. Early word embedding models like Word2Vec (Mikolov et al., 2013) and GloVe (Pennington et al., 2014) lacked contextual understanding. FastText (Bojanowski et al., 2017) built upon them and improved performance by incorporating subword information. Transformer (Vaswani et al., 2023) models like BERT (Devlin et al., 2019) and GPT (Radford et al., 2018) then revolutionized the field with context-aware embeddings. Multilingual models like mBERT, XLM (Lample and Conneau, 2019), and XLM-RoBERTa (Conneau et al., 2020) further advanced cross-lingual understanding, with recent open-source LLMs pushing performance even higher (Llama Team, 2024; Mistral AI, 2025). Using such Transformer models, documents and representative samples can be mapped into a shared embedding space to estimate their similarity. Focusing on transparency, simplicity and efficiency in our work, we use FastText and XLM-RoBERTa for our model-based filtering.
@@
Multilingual Evaluation. Evaluating LLMs requires diverse benchmarks testing linguistic and cognitive abilities like reading comprehension, reasoning, and knowledge. While established benchmarks such as MMLU (Hendrycks et al., 2020) and ARC (Clark et al., 2018) exist for English evaluation, assessments in other languages often rely on translations from English sources, as seen in XNLI (Conneau et al., 2018) and the machine-translated version of MMLU (Lai et al., 2023). However, translations can be problematic, failing to capture cultural nuances or introducing "translationese" (Romanou et al., 2024). Recent work by Romanou et al. (2024) and Singh et al. (2024a) emphasizes the importance of culturally sensitive, natively collected benchmarks. Task difficulty and formulation also impact model performance when trained for shorter durations (Kydlíček et al., 2024). In our work, we follow FineTasks, a recent evaluation tasks suite by Kydlíček et al. (2024) to assess our model-based filtering approaches across multiple languages.
@@ 3.1 Classifier Training Dataset
Aya Collection. A prompt completion dataset comprising 514M samples covering a variety of tasks, generated using instruction-style templates in 101 languages (Singh et al., 2024b).
@@
Include-Base-44. Multiple-choice questions focused on general and regional knowledge, extracted from academic and professional exams. Spanning 44 languages, it includes a total of 23K samples (Romanou et al., 2024).
@@
To address the quality difference, we create two Multilingual Knowledge Collection (MKC) configurations which allow us to evaluate the trade-off between data quality and scale:
@@
Dataset Creation. For our model-based filtering approaches, our goal is to identify documents from the pretraining dataset that are most similar to our representative samples, with the notion of similarity determined by the specific classifier used. We can directly measure similarity to our training data, for example, by calculating cosine similarity with training samples in the embedding space. Alternatively, following the approach of Li et al. (2024b), the task can be framed as a binary classification problem, with the representative samples as the positive class. For the negative class, we can subsample documents from our pretraining dataset, under the assumption that the majority of these documents are not well-structured or knowledge-rich. We use both approaches for our classifiers.
To create the binary classification training dataset, we selected up to 80K positive samples by using all examples from the smaller source datasets (e.g., Include-Base-44) and randomly subsampling from the Aya Collection for MKC+. The positive samples were combined with the same number of randomly sampled negative examples from FineWeb-2. The same training dataset was utilized across all model-based filtering approaches, disregarding negative samples when unnecessary. Additionally, we created a training dataset for each language individually to avoid leaking language-specific biases to data of other languages.
@@ 3.2 FastText-based Filtering (FT)
To efficiently process datasets with over 100 million documents (Penedo et al., 2024c), similar to DCLM (Li et al., 2024b), we used a binary FastText classifier (Joulin et al., 2017). FastText runs on CPU and can be deployed across multiple cores, for example using DataTrove (Penedo et al., 2024a).
We trained our FastText classifier on the processed training set using 2-gram features (4-gram for Chinese). These classifiers were then used to assign scores to all documents in the pretraining dataset. To filter the dataset, we applied a score threshold based on the desired retention percentage of documents. This approach balances dataset size and the predicted quality of the samples.
@@ 3.3 Transformer Embedding-based Filtering
To leverage rich semantic information based on contextual relationships, we utilized Transformer model embeddings. Specifically, we selected a pretrained XLM-RoBERTa base model (Conneau et al., 2020) due to its support of 100 languages, a relatively small size of 279M parameters, and its transparent training procedure. This choice enabled us to process web-scale data efficiently without being restricted to a single language and aligned with our commitment to open science.
@@
Multi-Layer Perceptron (MLP). We trained a single-hidden-layer neural network with a dimension of 256, the ReLU activation function, a 20% dropout, and the sigmoid function on the output. The network was trained for 6 epochs using the AdamW optimizer (Loshchilov and Hutter, 2019) with a constant learning rate and binary cross-entropy loss. We computed document scores using the output layer of the MLP model, which used XML-RoBERTa document embeddings as input.
@@ 4.1 Experimental Setup
Technical Details. We evaluate 1B-parameter Llama models (Llama Team, 2024) to demonstrate the effectiveness of our model-based filtering approaches. The models are trained on either 70B or 119B tokens, balancing token quality and diversity. The smaller dataset (70B tokens) exposes the model to each token at most once (with a few exceptions where some tokens appear twice). The larger dataset (119B tokens) simulates longer training, resulting in increased token repetition. Training utilizes the HuggingFace Nanotron library (Hugging Face, 2024a) with the AdamW optimizer (Loshchilov and Hutter, 2019) and a WSD learning rate schedule (Hägele et al., 2024).
@@
To minimize the need for costly hyperparameter tuning, we maintain a consistent setup across all experiments. Specifically, we adopt the DeepSeek scaling law (DeepSeek-AI et al., 2024) with a batch size of 1.6M tokens, learning rate of 0.0008, and 2000 warmup steps.
@@
As the base dataset, we use FineWeb-2 (Penedo et al., 2024c), which has been shown to provide a strong baseline across a variety of languages. Since FineWeb-2 is globally deduplicated, we rehydrate both filtered and unfiltered data using the hyperparameters recommended by Penedo et al. (2024c).
@@
To validate our method on English, we use three datasets: FineWeb (Penedo et al., 2024b) as the baseline, along with FineWeb-Edu (Penedo et al., 2024b) and DCLM (Li et al., 2024b), both of which represent the current state-of-the-art. Tokenization is performed using the multilingual Mistral v3 (Tekken) tokenizer (Mistral AI, 2024).
@@
Evaluation. Our evaluation prioritizes a diverse range of tasks to ensure the models retain well-rounded capabilities, rather than focusing exclusively on knowledge-based tasks. Specifically, we include tasks covering reading comprehension, general knowledge, natural language understanding, common-sense reasoning, and generative tasks in the target language. To evaluate our approach, we use the HuggingFace LightEval library (Fourrier et al., 2024).
@@
For French, Chinese, and Arabic, we utilize the FineTasks (Kydlíček et al., 2024) multilingual evaluation suite, which is designed to provide meaningful signals even for models trained in the order of 100B tokens. We select analogous tasks for German and Danish. For English, we rely on the SmolLM tasks suite (Hugging Face, 2024b). A complete list of tasks and their evaluation metrics for each language is provided in Appendix D.
@@
Model Selection. We follow the approach used in FineTasks (Kydlíček et al., 2024) for filter selection, computing a global rank score across individual metrics and languages to determine the optimal approach. For a detailed description of the average rank computation, please refer to Appendix E.
Computational Cost. We run our experiments on a compute cluster containing four GH200 chips per node, with each GH200 chip containing 72 CPU cores and one H100 GPU. Model training on 119B tokens costs approximately 1.1K H100 compute hours, while the embedding computation of all data for 20 languages costs approximately 4K H100 compute hours. We release the embeddings publicly so this computation does not have to be repeated666huggingface.co/datasets/epfml/FineWeb2-embedded. In total, we use approximately 152K H100 compute hours for running our experiments and generating the document embeddings. Data selection classifier training (i.e. FastText and MLP) takes a few minutes on a CPU. Filtering using any of the approaches (i.e., FastText, MLP, or CS) is computationally inexpensive, parallelizable, and is run on CPU. Overall, filtering German, Chinese, French, Arabic, and Danish data for one filtering approach in the ablations costs approximately 60 CPU hours, which is distributed over multiple CPU cores depending on the dataset size in each language to have filtering finish in approximately 30 minutes.